OpenStudy (anonymous):

What are the solutions of the quadratic equation? 4x2 + 34x + 60 = 0 ◘ 6, 2/3 ◘ –5/2 , –2/3 ◘ –6, –5/2 ◘ –6, 4

OpenStudy (anonymous):

That is x^2

OpenStudy (anonymous):

ok whats 1st?

OpenStudy (anonymous):

Well it is already equal to 0 , so you factor it?

OpenStudy (anonymous):

All the terms are even. You can divide everything by 2 first.

OpenStudy (anonymous):

Then, yes, it's pretty easy to factor.

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

do you know this equation?

OpenStudy (anonymous):

factoring : 4x^2+34x + 60 4x^2 +24x+10x+60 4x(x+6) + 10(x+6) (4x + 10) (x+6) = 0 can you do it now ?

OpenStudy (anonymous):

Yea, Haven't seen it since Algebra I though.

OpenStudy (anonymous):

a = 4 b = 34 c = 60 plug in and you will get the solution

OpenStudy (anonymous):

Okay, I think I got it.. Thank you.

OpenStudy (anonymous):

Woops, was supposed to be 2x^2 +17x +30 = 2x^2 +12x +5x +30 Then you can factor. (I think quadratic formula is overkill on this one...)